the primary care provider prescribes Flucanazole (Diflucan) 5mg/kg/day PO to be given in 2 equal doese. The client weighs 86 pounds. Available is fluconazole 40 mg/mL. How many mL will the nurse administer at each dose:
Answer #1

Ordered Flucanazole (Diflucan)= 5mg/kg/day PO to be given in 2 equal doese.  

Weight of client= 86 pounds= 86/2.2 kg= 39.09 kg

Dose as per the client;s weight = 5x 39.09= 195.4 mg per day to be given in 2 equal doses

means per dose amount= 195.4 /2= 97.72 mg

Availabele fluconazole = 40 mg/mL

Then required ml of fluconazole the nurse should administer at each dose will be = 97.72/40

=2.44 mL